# Case Law Grounding: Using Precedents to Align Decision-Making for Humans and AI

### v3

@@ 3. Case Law Grounding
In this work, we introduce and explore the framework of Case Law Grounding (CLG), a novel approach to supporting decision-making over subjective and socially-situated decisions aligned to the norms and preferences of groups and communities. We first examine concepts from the legal realm of case law and note key aspects that we adapt to create CLG. Then, we provide high-level details around the process for general decision-making following the CLG framework. Finally, we detail how CLG could be applied in two different settings: a human-led decision process for communities and an automated process that applies CLG with general-purpose large language models (LLMs).
3.1. Basis in Concepts from Legal Case Law
In judicial systems around the world, high stakes judgments are made regularly that depend on interpreting and applying concepts defined by society. In order to make consistent judgments on cases, some judicial systems have adopted the concept of “case law”, where in addition to higher-level statutory laws, past court decisions on substantially similar cases can also justify a position on a case.
We observe that case law systems rely on two main features to ensure the consistency and legitimacy of judgments without an exhaustive set of laws:
  • “Stand by what has been decided” (or Stare Decisis): Having a consistent criteria is crucial to any dependable system of making judgments—people want similar cases to have similar outcomes in order to perceive the system as fair. Case law systems promote stable judgments by encoding decision criteria implicitly into precedents. Rather than reason about the applicability of abstract criteria, which can vary widely due to subjective interpretations, judges can instead reason about the more constrained question of whether a case is similar to this or that decided case.
  • “Trust but verify” through Legal Citations: Making judgments often involves choosing between opposing options which naturally result in disagreements. When drafting a decision in the legal space, lawyers and judges provide citations to prior cases not just as context but in order to situate the case in the broader graph of dependencies between cases. In this way, the validity of cited cases provides an attestation to the soundness of a decision. By maintaining proper citations, the rationale for decisions can be inspected by anyone through following citation chains.
However, because of the high-stakes nature of cases in the legal system, judgments are typically conducted by experts with years of training and carried out in formal settings where significant time is allotted for deliberation. We must make adaptations to traditional case law before we can apply it to decision-making tasks such as content moderation or alignment of AI models that have significantly higher throughput but less existing structural support (such as detailed reasoning around case judgments).
@@ 3.1. Basis in Concepts from Legal Case Law
A diagram of the case law grounding process implemented both as
3.2. Using Cases to Ground Decision-Making
3.2.1. Step 1: Retrieving Relevant Cases
@@
The first step in case law involves conducting legal research. In the legal practice, lawyers spend considerable effort to locate, examine, and reason about the space of past decided cases in order to draft the best argument for their client. This often involves querying legal databases to identify relevant cases that can be binding precedents for the case they are arguing (Berring, 1987). Once relevant cases are found, there can be additional criteria and checks to make sure that it can apply.
@@
Semantic Similarity for Case Retrieval: Experts in the legal space depend on specialized keywords and indexing systems that let them navigate well reasoned facets of case law to quickly localize their search space. These indices are valuable because they can encode crucial aspects of a case that directly help identify its viability as a precedent (e.g., is the jurisdiction the same, are there commonalities in the circumstances, evidence, or statutory nature of the case); thus it makes sense to spend resources to create them. In the settings we consider, we can’t expect to have ready-made case indices, but we can utilize advances in general purpose language models to approximate case retrieval through text embeddings. Since the cases we’re interested in (e.g., online comments, interpersonal advice scenarios, user queries, and AI model responses) encode all of their factual properties through their natural language descriptions, a reasonable proxy for the relevance of any case to any other can be approximated in the form of the distance between them in the embedding space $d(x_{\text{past}},x_{\text{current}})$. Depending on the domain, it may be possible to choose an embedding specifically created for the domain, though in many situations, an off-the-shelf generic embedding can also be sufficient.
3.2.2. Step 2: Determining Applicable Precedents
@@
Simply surfacing potentially relevant cases to the agent making a decision is likely not sufficient, as we don’t know whether the decision-maker considers the prior case as precedent setting and whether they have integrated it into their judgment. In the legal realm, the selection of precedents from related cases is made explicit, taking the form of a written document containing the arguments being made, and especially the citations to prior precedent cases that support such arguments. Indeed, this explicit identification and documentation of precedents as supporting arguments to a judgment can confer many benefits. For one, by documenting the exact cases used to support a judgment, it becomes easier to audit the reasoning behind judgments. One can gain understanding of disagreements between different decision-makers by inspecting the commonalities and differences in the precedents selected. Additionally, selecting precedents decouples the judgment process from a more subjective (and possibly un-grounded) interpretation of a case and outcome to instead focus on reasoning about more objective comparisons between aspects of the cases. Finally, by having a shared set of precedents to draw from, we can improve consistency of decisions across different adjudicators who may have different personal subjective preferences.
@@
However, requiring a full argument for each case, while effective for challenging crowd judgment tasks (Drapeau et al., 2016; Chen et al., 2019), can be labor intensive. To mitigate this issue, in this stage we simplify the argument process to focus only on classification of whether a retrieved case is a good precedent, not why it is so. This means that instead of coming up with a judgment, writing a justification, and citing cases to support it, an adjudicating agent focuses on presenting their evaluation of the neighborhood of cases.
@@
  • Precedents: A retrieved case should be marked as a precedent if that prior case provides support for the judged case to adopt a similar decision due to substantially similar circumstances or aspects (i.e., cases that fall on the same side of the decision bound of the judged case).
  • Doesn’t Apply: A retrieved case does not apply as a precedent if there are critical aspects of the case that differ from the judged case such that the argument for the past decision would not apply to the new case. While many cases that don’t apply as precedents are likely to fall on the opposite side of the decision bound of the judged case, this is not a requirement for the case to not apply. Sufficiently significant differences in circumstance or factual aspects don’t necessarily need to result in opposing outcomes (e.g., involvement of a minor versus adult in a legal case might affect the argument but the outcome judgment could end up being the same).
@@
3.2.3. Step 3: Deriving Judgments from Selected Precedents
@@ 3.3. Supporting Human-Led and LLM Prompted Decision-Making
3.3.1. Case Selection Tool for Human-Led Decision-Making
@@
We use the right side interface to facilitate the steps for conducting decision-making following or case law grounding process. During annotation, the right side of the interface is populated with a list of cases retrieved through the case retrieval step (3(b)). Under each retrieved case, the annotator can click on a button to toggle whether a case should be considered a precedent or not. Once the annotator is done evaluating each case, they are asked to provide a final decision given their review of the precedents but not constrained to the decisions of them. While selecting the set of applicable precedents alone is sufficient for the CLG workflow, this last question is included to allow us to evaluate whether enforcing binding precedents (subsubsection 3.2.3) is important to final accuracy or not.
@@ 3.3.2. LLM Prompts for Precedent Selection
To evaluate an automated process where LLMs contribute to decision-making instead of human annotators, we also implemented a series of prompts for LLMs to follow a grounding process similar to the human-led version (Appendix A). In the prompting version of the workflow, we set up the system prompt in a similar way to the instructions section in the annotation interface, providing the instructions for the decision task itself (whether the model is assessing case precedents or constitutional guidelines) as well as task domain specific instructions. We then provide the case being judged along with the grounding support as part of the user prompt. Each user prompt assessment was prompted independently, with no conversation history persisted between two different user prompt queries.
For the LLM-prompting version of CLG, we use a case selection prompt (subsection A.2) to evaluate each retrieved case, substituting {input} with the case being judged, {precedent} with one retrieved past case, and {decision} with its past decision. This prompt was then repeatedly invoked for each retrieved case, producing a binary output for whether the case should be considered a precedent or not. Following the CLG process, these precedent determinations were then used to derive the final decision. Unlike the human annotation interface where annotators also produced a final decision based on their review of precedents but unconstrained by the decisions of those precedents, we could not create a prompt that provided all precedents to the model to produce a single final decision in a way that fit within the context window of the model, thus non-binding precedents were only evaluated for the human-led process and not the LLM prompting version. For the LLM-prompting version of constitutional grounding, we use a guideline assessment prompt (subsection A.3) to evaluate each retrieved case, substituting {input} with the case being judged and {instructions} with a list of rules and guidelines for making the decision. We then collected the resulting output directly as the model’s final decision.
@@ 4. Experiments
As the goal of CLG is to support decision-making on subjective and socially-situated tasks for groups and communities, we focused our evaluation of CLG on these types of tasks, drawing from two common subjective decision tasks: content moderation and toxicity rating. We then focused on sampling several distinct groups or communities which have been observed by prior work to have distinct preferences and norms when it comes to making decisions on these subjective tasks to compare decisions made with support from traditional constitutional grounding (rule) against decisions made with case law grounding (case) support. We set out to answer the following research questions in settings where humans are the decision-makers as well as settings where LLMs are being prompted to make decisions:
  • RQ1: For subjective and socially-situated decisions, does the case law grounding approach result in more accurately aligned decisions compared to traditional rule-based constitutional grounding?
  • RQ2: For case law grounding, how do different choices in the retrieval setup and precedent selection affect the alignment accuracy of decisions produced?
@@
    • RQ2b: How does enforcing binding precedents affect the alignment accuracy of the decisions?
  • RQ3: For the automated LLM-prompting based process, do our findings around CLG continue to scale as we move to larger datasets / additional population groups?
4.1. Experiment Setup
For RQ1, we set up our experiments to focus on evaluating how well the decisions produced by different grounding approaches aligned with the ground truth, comparing two main conditions rule—which represents a constitutional grounding process, and case—which represents the case law grounding approach in CLG. To more comprehensively evaluate these two conditions, we conducted this comparison across 5 groups or communities each of which were making decisions on one of 2 task domains. In addition to groups, we also evaluated whether our findings between conditions holds across different versions of the process that used different agents for making the decisions: human-led and llm-prompting based. For RQ2, we examine how different configurations for CLG may affect performance, with RQ2a assessing different retrieval window sizes via simulation and RQ2b comparing an additional condition example for the human-led process that represents a weaker form of case law grounding. Finally, for RQ3, we only look at the LLM-prompting version of both CLG and constitutional grounding, comparing whether they retain their relative performance on a additional sets of cases that we did not collect human annotations for.
4.1.1. Tasks and Datasets
We picked two real-world decision-making tasks that involve subjective decisions with social context specific to groups: a comment moderation task where the decision is to “keep” or “remove” a comment posted in a Reddit community, and a toxicity rating tasks where the decision is the scalar rating of the toxicity level of a social media post (on a scale from 1–5). Then, for each task, we selected a dataset that contained ground truth annotations provided by: (1) different groups of users, (2) that we could recruit for our annotation study.
For the mod domain, we used the NormVio dataset (Park et al., 2021) which contains comments from various communities on Reddit that may or may not have violated the norms of that community and as a result have been evaluated by moderators for potential removal. This dataset separates comments by the “subreddit” it was posted in, which we use to partition distinct communities that will likely have distinct preferences around content moderation. From this dataset, we selected the top 10 subreddits with a sufficiently large size of examples ($n>100$) and did not exhibit significant skew towards either decisions of “keep” or “remove”. We then further narrowed down this set based on avoiding multiple communities around similar topics and prioritizing communities that were easier to recruit from, ultimately sub-sampling this dataset down to 3 communities: r/Coronavirus, r/AmItheAsshole, r/Gaming. Ground truth labels for each case is established by the actual moderation decision applied to each comment.
For the tox domain, we used the toxicity perspectives dataset (Kumar et al., 2021) which contains a wide range of posts collected from social media platforms that were then evaluated by crowd annotators and rated for toxicity. This dataset contains disaggregated ratings that are associated with demographic information of the annotator, which we use to segment a general crowd worker population into distinct salient groups. Based on the findings from the original paper, we sampled from 3 disjoint intersectional demographic slices (see subsubsection 4.1.2) that were indicated to have the most significant differences with each other: seen=N,targeted=N, seen=Y,targeted=N, seen=Y,targeted=Y. Ground truth labels for each case in each group is established by taking the mean rating for that group rounded to the nearest integer.
For each slice in each task domain above, we first shuffled the instances randomly, then drew 10% of cases only used as past precedents, 10% judged by both the human-led process and LLM prompting (rounded down to nearest multiple of 10 to facilitate annotation), and 10% only evaluated with LLM prompting. To ensure a consistent level of effort for the annotation task, cases set to be judged by human annotators were partitioned into batches of 10, with each annotator only judging one such batch within their session. When evaluating the constitutional grounding condition (rule), we use published community guidelines for each subreddit as approximations of a high-quality constitution for groups in mod. Meanwhile, for the tox domain, we adapt the specific instructions and criteria given to annotators for rating toxicity in the original dataset from Kumar et. al. as a list of rules that serve as the constitution for all groups. Finally, to address the fact that community guidelines may have changed over time, we spot checked the community guidelines for each of the 3 groups in the mod domain using the Wayback Machine [^1] https://web.archive.org/ to confirm that no significant changes (such as adding or removing rules) were made to the guidelines across the time span of cases evaluated.
4.1.2. Communities and Groups
One challenge of decision-making around socially-constructed criteria so challenging is the abundance of localized preferences over criteria that differ across different population groups or communities. As CLG is intended to support making decisions around subjective concepts in a way that aligns with the norms and preferences of a group, we set up our evaluation to include two types of social groupings that are likely to have consistent norms and preferences around subjective decisions: interest-based communities—consisting of participants recruited from active users within a subreddit, and demographic-based population groups—consisting of general participants that report belonging to some demographic group.
For the task of content moderation (mod), where decisions are made around whether content is removed or retained, prior work has observed that different communities can often have significantly different preferences when it comes to adhering to a formally defined criteria or values (Koshy et al., 2023). Thus, drawing from our dataset, we selected 3 communities from Reddit: r/Games—a broad interest group around game content and the games industry, r/Coronavirus—a community focused around high-quality science-informed discussions related to COVID-19, and r/AmItheAsshole—a collective space for sharing and opining over a variety of interpersonal conflict scenarios.
For the task of toxicity rating (toxicity), prior work has found that differences in the background of raters, such as their prior exposure to toxic content as observers or victims, can affect how they rate the same piece of text (Kumar et al., 2021). While raters do not necessarily form a “community”, we can still view demographic features as a way to separate the population into groups that represent distinct norms and preferences. In our case, we stratify the general population by taking two demographic properties identified to be significant correlating factors (Kumar et al., 2021) of toxicity ratings: whether a rater has often seen toxic content when interacting with social media (negatively correlated with scores), and whether the rater has been personally targeted by toxic content (positively correlated with scores). Noting that it does not make sense to have been targeted yet never seen toxic content, this results in 3 final groups: seen=Y,targeted=N, seen=N,targeted=N, and seen=Y,targeted=Y. During recruitment, we found it challenging to find participants for seen=N,targeted=N, so we only evaluated the LLM-prompting process (RQ3) with that group, excluding it from other evaluations.
4.1.3. Participant Recruitment and Crowdsourced Annotation
As we envision human-led processes for grounding subjective decisions (whether constitutional or case law) to be used in practice by groups and communities themselves rather than by external populations, we set out to recruit participants that would demographically best match the original group or community producing the ground truth data. To achieve this goal, we applied pre-screening surveys in each of our recruitment efforts (Appendix B), with participants first asked to fill out the screening survey, after which we would select and assign tasks conditions to them.
As each of the mod groups represents a specific subreddit community, we recruited participants by running Reddit Ads targeted to subreddits used in our task domains. Participants were asked to fill out the screening survey to self report subreddits they frequented within a list of communities (subset of the original top-10 candidates subsubsection 4.1.1, removing 1 subreddit that had been removed). Then, participants who reported frequenting at least one of the 3 target subreddits would be allocated a batch and condition, balancing such that each community, condition, and batch had a target of 3 annotators. At the end, this resulted in annotations on 3 batches for r/AmItheAsshole, 2 for r/Games, and 1 for r/Coronavirus.
In the tox domain, original annotations were provided by Amazon Mechanical Turk (AMT) workers. However, recent work has shown concerns around AMT (Marshall et al., 2023) annotations that have developed since the original dataset was collected. To avoid issues with introducing significantly more quality control components than the original task, we opted to recruit from Prolific, as it is a similar crowd work platform with (as of our deployment) lower rates of data quality issues. Participants were screened with two questions about their prior exposure to toxic posts with these answers then used to assign them to one of 3 population groups (see 4.1.2). During recruitment, we also found that no crowd workers reported being in the seen=N,targeted=N group during screening. While this was ultimately unsurprising, we opted to drop this group from the evaluations related to human-led processes (RQ1 and RQ2), and only used this group for the final LLM experiments around scaling.
As both experiment domains involve reading content that is potentially offensive, we made sure to prominently disclose this aspect during recruitment. We also allowed participants to drop out during the task, if they experienced discomfort with the annotation. Human participants in both conditions were paid a base rate of \$15 for completing our annotation study (30-45 minutes), with Prolific participants being paid through the platform and Redditors being paid in the form of an Amazon Gift Card. Annotators electing to withdraw consent during the task were pro-rated payment for the amount of annotations completed, with their data excluded from the evaluation. Additionally, to incentivize participation from Redditors and reduce drop-out between screening and completion, we provided an additional bonus of \$5 to Redditors who completed our annotation task within 24 hours of assignment. In total, we were able to recruit $n=60$ participants from Prolific and $n=35$ participants from Reddit.
4.1.4. Additional Setup Details
4.2. RQ1: Comparing Case Law and Constitutional Grounding

### v4

@@ 3. Case Law Grounding
In the legal space, there are similar challenges when it comes to grounding decisions in a way that follows the will of the public, and prior work has observed similarities between the grounding of decisions in sociotechnical settings and those in the legal realm (Caputo, 2024). One interesting parallel lies in the mechanisms used for grounding decisions: similar to the rules and guidelines common in sociotechnical systems, constitutional (or statutory) laws also serve a similar purpose of providing high-level guidance around decisions. However, in addition to constitutional grounding, some legal systems also make use of case law to address situations when constitutional guidance is ambiguous, incomplete, or conflicting. The concept of case law presents the idea that past decisions can also ground future decisions in the form of analogies between past cases and the current case. By identifying similarities and differences in the nature of facts within each case, it is possible to derive ad-hoc unstated principles or determinations that people are in agreement on (Ashley, 1991).
In this work, we propose Case Law Grounding (CLG) as an initial approach that explores the idea of using past decisions to ground decision-making (Figure 1). In the following subsections, we will describe a general set of steps for conducting case law grounding in the decision-making process, followed by two specific implementations of this process in a human-led and an LLM-prompted design.
@@
A diagram of the case law grounding process implemented
3.1. Step 1: Retrieving Relevant Cases
@@
The first step in case law involves conducting legal research. In legal practice, lawyers spend considerable effort to locate, examine, and reason about the space of past decided cases in order to draft the best argument for their client. This often involves querying legal databases to identify relevant cases that can be binding precedents for the case they are arguing (Berring, 1987). Once relevant cases are found, there can be additional criteria and checks to make sure that they can apply.
@@
Semantic Similarity for Case Retrieval: Experts in the legal space depend on specialized keywords and indexing systems that let them navigate well-reasoned facets of case law to quickly localize their search space. These indices are valuable because they can encode crucial aspects of a case that directly help identify its viability as a precedent (e.g., is the jurisdiction the same, are there commonalities in the circumstances, evidence, or statutory nature of the case); thus it makes sense to spend resources to create them. In the settings we consider, we can’t expect to have ready-made case indices, but we can utilize advances in general-purpose language models to approximate case retrieval through text embeddings. Since the cases we’re interested in (e.g., online comments, interpersonal advice scenarios, user queries, and AI model responses) encode all of their factual properties through their natural language descriptions, a reasonable proxy for the relevance of any case to any other can be approximated in the form of the distance between them in the embedding space $d(x_{\text{past}},x_{\text{current}})$. Depending on the domain, it may be possible to choose an embedding specifically created for the domain, though in many situations, an off-the-shelf generic embedding can also be sufficient.
3.2. Step 2: Determining Applicable Precedents
@@
Simply surfacing potentially relevant cases to the agent making a decision is likely not sufficient, as we don’t know whether the decision-maker considers the prior case as precedent setting and whether they have integrated it into their judgment. In the legal realm, the selection of precedents from related cases is made explicit, taking the form of a written document containing the arguments being made, and especially the citations to prior precedent cases that support such arguments. Indeed, this explicit identification and documentation of precedents as supporting arguments to a judgment can confer many benefits. For one, by documenting the exact cases used to support a judgment, it becomes easier to audit the reasoning behind judgments. One can gain an understanding of disagreements between different decision-makers by inspecting the commonalities and differences in the precedents selected. Additionally, selecting precedents decouples the judgment process from a more subjective (and possibly un-grounded) interpretation of a case and outcome to instead focus on reasoning about more objective comparisons between aspects of the cases. Finally, by having a shared set of precedents to draw from, we can improve the consistency of decisions across different adjudicators who may have different personal subjective preferences.
@@
However, requiring a full argument for each case, while effective for challenging crowd judgment tasks (Drapeau et al., 2016; Chen et al., 2019), can be labor intensive. To mitigate this issue, in this stage we simplify the argument process to focus only on the classification of whether a retrieved case is a good precedent, not why it is so. This means that instead of coming up with a judgment, writing a justification, and citing cases to support it, an adjudicating agent focuses on presenting their evaluation of the neighborhood of cases.
@@
  • Precedent [^1] Often referred to as an “adopted” precedent.: A retrieved case should be marked as a precedent if that prior case provides support for the judged case to adopt a similar decision due to substantially similar circumstances or aspects (i.e., cases that fall on the same side of the decision bound of the judged case).
  • Doesn’t Apply [^2] Often referred to as a “distinguished” precedent.: A retrieved case does not apply as a precedent if there are critical aspects of the case that differ from the judged case such that the argument for the past decision would not apply to the new case. While many cases that don’t apply as precedents are likely to fall on the opposite side of the decision bound of the judged case, this is not a requirement for the case to not apply. Sufficiently significant differences in circumstance or factual aspects don’t necessarily need to result in opposing outcomes (e.g., involvement of a minor versus adult in a legal case might affect the argument but the outcome judgment could end up being the same).
@@
3.3. Step 3: Deriving Judgments from Selected Precedents
@@ 3.4. Supporting Human-Led and LLM Prompted Decision-Making
3.4.1. Case Selection Tool
@@
We use the right side interface to facilitate the steps for conducting decision-making following the case law grounding process. During annotation, the right side of the interface is populated with a list of cases retrieved through the case retrieval step (3(b)). Under each retrieved case, the annotator can click on a button to toggle whether a case should be considered a precedent or not. Once the annotator is done evaluating each case, they are asked to provide a final decision given their review of the precedents but not constrained to the decisions of them. While selecting the set of applicable precedents alone is sufficient for the CLG workflow, this last question is included to allow us to evaluate whether enforcing binding precedents (subsection 3.3) is important to final accuracy or not.
@@ 3.4.2. LLM Prompts for Precedent Selection
To evaluate an automated process where LLMs contribute to decision-making instead of human annotators, we also implemented a series of prompts for LLMs to follow a grounding process similar to the human-led version (Appendix B). In the prompting version of the workflow, we set up the system prompt in a similar way to the instructions section in the annotation interface, providing the instructions for the decision task itself (whether the model is assessing case precedents or constitutional guidelines) as well as task domain specific instructions. We then provide the case being judged along with the grounding support as part of the user prompt. Each user prompt assessment was prompted independently, with no conversation history persisted between two different user prompt queries.
For the LLM-prompting version of CLG, we use a case selection prompt (subsection B.2) to evaluate each retrieved case, substituting {input} with the case being judged, {precedent} with one retrieved past case, and {decision} with its past decision. This prompt was then repeatedly invoked for each retrieved case, producing a binary output for whether the case should be considered a precedent or not. Following the CLG process, these precedent determinations were then used to derive the final decision. Unlike the human annotation interface where annotators also produced a final decision based on their review of precedents but were unconstrained by the decisions of those precedents, we could not create a prompt that provided all precedents to the model to produce a single final decision in a way that fit within the context window of the model; thus, non-binding precedents were only evaluated for the human-led process and not the LLM prompting version. For the LLM-prompting version of constitutional grounding, we use a guideline assessment prompt (subsection B.3) to evaluate each retrieved case, substituting {input} with the case being judged and {instructions} with a list of rules and guidelines for making the decision. We then collected the resulting output directly as the model’s final decision.
@@ 4. Experiments
To evaluate CLG, we set out to answer the following research questions:
  • RQ1: Does the case law grounding approach result in more accurately aligned decisions compared to traditional rule-based constitutional grounding?
  • RQ2: How do specific choices in the retrieval setup and precedent application affect the alignment accuracy of decisions produced?
@@
    • RQ2b: How do binding precedents affect the alignment accuracy of the decisions?
@@
4.1. Experiment Setup
To answer RQ1, we set up two main conditions rule—representing a constitutional grounding process, and case—representing CLG. We conducted this comparison across 2 types of decision-making tasks (moderation and toxicity rating) spanning 5 community / group specific decision-making task instances. In each condition, we evaluated both the human-led and llm-prompting versions of CLG.
For RQ2, we dive into how the choices in the CLG configuration contributed to performance by conducting ablation-style experiments. Specifically, we examined the effects of a smaller retrieval window size on decision accuracy (RQ2a) and we compared the decisions derived from binding precedents with human-produced decisions (non-binding) to examine the need for binding precedents (RQ2b).
4.1.1. Tasks and Datasets
We picked two real-world decision-making tasks that involve subjective decisions with social context specific to groups: a comment moderation task where the decision is to “keep” or “remove” a comment posted in a Reddit community, and a toxicity rating task where the decision is the scalar rating of the toxicity level of a social media post (on a scale from 1–5). We selected datasets for ground truth annotations that covered different groups / communities that we could readily recruit for our annotation study. This dataset separates comments by the “subreddit” it was posted in, which we use to partition distinct communities that will likely have distinct preferences around content moderation. For the tox domain, we used the toxicity perspectives dataset (Kumar et al., 2021) which contains a wide range of posts collected from social media platforms that were then evaluated by crowd annotators and rated for toxicity. This dataset contains disaggregated ratings that are associated with demographic information of the annotator, which we use to segment a general crowd worker population into distinct salient groups.
Within each dataset, we performed additional steps to select for communities and groups with sufficient data size that we could readily recruit that we document in subsection A.1. We also randomly shuffled and sub-sampled data for each task domain and group instance (10% as precedents, 10% used in both human and LLM decision evaluation). We also divided sampled cases into “batches” of 10 each, with each annotator only judging one such batch. In the constitutional grounding condition (rule) we drew from existing rules and guidelines associated with each task domain and group (see subsection A.1 for details).